OpenStudy (anonymous):

Algebra A rectangle of area 350 square feet is 14 times as wide as it is long. Find it's Length and Widith? Please Show Work (: Algebra A rectangle of area 350 square feet is 14 times as wide as it is long. Find it's Length and Widith? Please Show Work (: @Mathematics

OpenStudy (anonymous):

Is that my whole problem and answer?

OpenStudy (anonymous):

\[\huge 350 feet^2=xy\] \[\huge x=14y\Rightarrow14y^2=350feet^2→y=5feet\] \[\huge x=14*5feet=70feet\]